# On structural contraction of biological interaction networks

### v3

@@ 1 Introduction
Figure 1: The same BIN formalism $S+K\leftrightharpoons C\to P+K$ can represent processes as varied as: (a) an interaction between a catalyst (in the figure, an enzyme kinase) and its substrate, in which the two species can temporarily interact in a reversible fashion, sometimes resulting in the catalyst being released and the substrate being modified (“product”); (b) an interaction between an infective individual and a susceptible individual, in which the two individuals meet and either go their way without any new infection, or result in the susceptible individual becoming exposed to the disease; or (c) a predator/prey interaction in which the prey might not be harmed or the pray may end up injured or dead.
@@
We introduce the notion of structural contraction, and show that how this property can be deduced from the conditions for stability introduced in [4, 5] and implemented in the computational package LEARN. In particular, we show that the existence of the rate-dependent Lyapunov functions proposed in [4, 5] for a given reaction network implies structural nonexpansivity with respect to weighted $\ell_{\infty}$ norms, and strict contractivity on positive compact sets under additional but easily verifiable conditions. Furthermore, we show that such networks obey the notion of “contraction after a small overshoot and short transient” that was introduced in [63], and this fact allows us to show entrainment to periodic inputs. Therefore, the powerful tools of contraction theory can be applied to a rich class of BINs which can be characterized computationally via the MATLAB package LEARN (github.com/malirdwi/LEARN) or graphically [1].
The paper is organized as follows. We first review the main results. Section 2 discusses preliminaries including extensions of previous results that are relevant to our current analysis. Section 3 shows that nonexpansivity follows from the existence of an appropriate rate-dependent Lyapunov function. Section 4 establishes contraction over positive compact sets. Section 5 shows contraction after a small overshoot and short transient and entrainment to periodic inputs. Section 6 discusses biochemical examples in details, and conclusions are discussed in Section 7.
@@ 1.1 Overview of Results
In this subsection, we offer an informal guide that summarizes the main results of the paper. They apply to BINs whose dynamics can be described by $\dot{x}=\Gamma R(x)$ where $\Gamma$ is the stoichiometry matrix and $R$ is the reaction rate. The only assumption needed is that the kinetics are monotone (refer to §2 for the exact assumptions)
@@
The results in [4, 5] provided functions that can be defined using the stoichiometry of the network without knowledge of the kinetics. We first define a rate-dependent Graphical Lyapunov Function (GLF) ${\widetilde{V}}(r)$, where $r$ has the dimension equal to the number of reactions, see Definition 2.1. It is characterized by the existence of a common Lyapunov function for a set of rank-one linear systems, see Theorem 2.4. This means that $V(x)={\widetilde{V}}(R(x))$ is a Lyapunov function which is always non-increasing along the dynamics, see Proposition 2.5. In addition, if ${\widetilde{V}}$ is invariant with respect to translation by $\Gamma$, then its existence is equivalent to the existence of a dual function ${\widehat{V}}$ satisfying ${\widehat{V}}(\Gamma r)={\widetilde{V}}(r)$, see Theorem 2.12. Then, ${\widehat{V}}(x-{\bar{x}})$ is a Lyapunov function in the stoichiometric class compatible with a given steady state ${\bar{x}}$, see Corollary 2.13.
@@
Our main result states that if an $\ell_{\infty}$ GLF ${\widetilde{V}}(r)=\|B\Gamma r\|_{\infty}$ exists for a network, then the norm of the difference between two trajectories that lie in the same stoichiometry class is non-increasing along time, see Theorem 3.4, thus establishing that the network is structurally nonexpansive with respect to the norm .$\|.\|_{B}:z\mapsto\|Bz\|_{\infty}$. Furthermore, we provide an upper bound on the logarithmic norm of the Jacobian in terms of the partial derivatives of $R$, see Lemma 3.9.
@@
Given the satisfaction of these two verifiable conditions, we show that a network is structurally strictly contractive over positive compact sets with respect to the norm .$\|.\|_{PB}:z\mapsto\|PBz\|_{\infty}$, where $P$ is an appropriately-defined diagonal matrix that might depend on the choice of the compact set, see Theorem 4.10.
@@
We then turn to the study of BINs where one or multiple parameters are periodic time-varying functions. We ask if all trajectories entrain to the same period. To that end, we use our result establishing contraction over positive compact sets to show when a given BIN is also contractive with short transients and small overshoots, see Theorem 5.5. Once that this is known, the results from [63], allow us to conclude that trajectories of a periodic time-varying system that stay in the same stoichiometric class entrain to periodic inputs, see Theorem 5.7.
@@ 2.1 Notations
For any set $S$, $|S|$ denotes its cardinality. The set of $n$-tuples of positive numbers is denoted by $\mathbb{R}_{+}^{n}$, while the set of $n$-tuples of non-negative numbers is denoted by $\mathbb{R}_{\geq 0}^{n}$. The set of $n\times m$ matrices with real entries is denoted by $\mathbb{R}^{n\times m}$. For any subset $U\subset\mathbb{R}^{n\times m}$, its closure with respect to the standard Euclidean topology is denoted by $\bar{U}$. If $A\in\mathbb{R}^{n\times m}$ is a matrix, then ${\mathop{\rm Ker\,}}A$ is its kernel or null space, while ${\mathop{\rm Im\,}}A$ is its image space. For any function $V:\mathcal{X}\to\mathbb{R}$, ${\mathop{\rm Ker\,}}V:=\{x\in\mathcal{X}|V(x)=0\}$, and for a subset $\mathcal{Y}\subset\mathcal{X}$, ${\mathop{\rm Ker\,}}V|_{\mathcal{Y}}:=\{x\in\mathcal{Y}|V(x)=0\}$. For $x=[x_{1},..,x_{n}]^{T}\in\mathbb{R}^{n}$, $\|x\|_{\infty}:=\max_{i}|x_{i}|$. The notation $x\gg 0$ means that $x\in\mathbb{R}_{+}^{n}$. For a square matrix $A\in\mathbb{R}^{n\times n}$, $\sigma_{i}(A):=a_{ii}+\sum_{j\neq i}|a_{ij}|$, $i=1,..,n$, and $\mu_{\infty}(A):=\max_{i}\sigma_{i}(A)$. For a subspace $W\subset\mathbb{R}^{n}$, the quotient space $\mathbb{R}^{n}/W$ is the set of all sets of the form $\{x+y|y\in W\}$ where $x\in\mathbb{R}^{n}$.
@@ 2.2 Biological Interaction Networks
A Biological Interaction Network (BIN) $\mathscr{N}$ is specified as a pair $(\mathscr{S},\mathscr{R})$ where $\mathscr{S}:=\{X_{1},...,X_{n}\}$ is the set of symbols called species, and :=$\mathscr{R}:=\{{\rm\bf R}_{1},..,{\rm\bf R}_{\nu}\}$ is a set of symbols called reactions. Each reaction is associated to a formal rule of the following form:
${\rm\bf R}_{j}:~{}~{}\sum_{i=1}^{n}X_{i}\longrightarrow\sum_{i=1}^{n}\beta_{ij}X_{i},$
where the nonnegative real numbers $\alpha_{ij},\beta_{ij}\geq 0$ are called the stoichiometric coefficients of the respective reaction. The $X_{i}$’s that appear with nonzero coefficients in the left-hand side are called the reactants of the reaction $j$, and the $X_{i}$’s that appear with nonzero coefficients in the right-hand side are called the products of the reaction $j$. In many applications the stoichiometric coefficients are integers, and one thinks of a reaction as saying that $\alpha_{i1}$ copies of $X_{1}$, …, $\alpha_{in}$ copies of $X_{n}$ combine to produce $\beta_{i1}$ copies of $X_{1}$, …, $\beta_{in}$ copies of $X_{n}$. The net “loss” (if negative) or “gain” (if positive) of the $i$th species in the $j$th reaction is $\gamma_{ij}=\beta_{ij}-\alpha_{ij}$. These numbers can be collected into a matrix $\Gamma\in\mathbb{R}^{n\times\nu}$ defined entry-wise as $[\Gamma]_{ij}=\gamma_{ij}$. The matrix $\Gamma$ is customarily called the stoichiometry matrix.
@@
In order to quantify elements of the network, the species $X_{1},..,X_{n}$ are assigned non-negative numbers known as the concentrations (representing the density of elements in a certain region of space) $x_{1},..,x_{n}$, while reactions ${\rm\bf R}_{1},..,{\rm\bf R}_{\nu}$ are assigned rates $R_{j}:\mathbb{R}_{\geq 0}^{n}\to\mathbb{R}_{\geq 0}^{\nu},j=1,..,\nu$. These rates describe “how fast” each reaction can occur, depending on the concentration of its reactants. The most common form of reaction rates is known as mass-action kinetics, which are those of the form $R_{j}(x)=\prod_{i=1,\alpha_{ij}>0}^{n}k_{j}x^{\alpha_{ij}}$, where $k_{j}>0$ is known as the kinetic constant. However, here we do not need to restrict to a specific form of kinetics. Instead we only assume that each such $R_{j}$ is monotone with respect to its reactants. More precisely, a reaction rate $R$ is admissible if it satisfies the following general properties:
@@
    each reaction varies smoothly with respects to its reactants, i.e $R(x)$ is $\mathscr{C}^{1}$;
@@
    if a reactant increases, then the reaction rate increases, i.e ${\partial R_{j}}/{\partial x_{i}}(x)\geq 0$ if $\alpha_{ij}>0$ and ${\partial R_{j}}/{\partial x_{i}}(x)\equiv 0$ if $\alpha_{ij}=0$. Furthermore, the aforementioned inequality is strict whenever the reactants are strictly positive.
@@
Another equivalent formulation is obtained by decomposing $R$ into a conic sum of indicator matrices. To that end, let $E_{ji}\in\{0,1\}^{\nu\times n}$ be an indicator matrix, i.e, $E_{ij}$ is zero everywhere except for the $(j,i)$th entry where it is equal to 1. In addition, let the elements of $\mathcal{P}_{\mathscr{N}}$ be indexed as $(i_{1},j_{1}),..,(i_{s},j_{s})$. Therefore, we can write
(2) $\mathcal{K}_{\mathscr{N}}=\{K\in\mathbb{R}^{\nu\times n}|K=\sum_{\ell=1}^{s}\bar{\rho}_{\ell}E_{j_{\ell}i_{\ell}},~{}\mbox{for some}~{}\bar{\rho}_{1},..,\bar{\rho}_{s}>0\}.$
Hence, the admissibility of $R$ implies the existence of continuous functions $\rho_{1}(x),..,\rho_{s}(x)\geq 0$ such that
@@
The time evolution of the concentration vector $x=[x_{1},..,x_{n}]^{T}$ as a function of time is by definition described by the following vector Ordinary Differential Equation (ODE):
@@
[…]
@@
The forward-invariant subset $\mathscr{C}_{x_{0}}$ is a closed and convex set known as the stoichiometric class corresponding to $x_{0}$.
@@
[…]
@@
$\dot{x}(t)\;=\;\Gamma\dot{\xi}(t)\;=\;\Gamma R(x_{0}+\Gamma\xi(\tau))\,d\tau\;=\;\Gamma R(x(\tau))\,d\tau\,.$
@@ 2.3 Graphical Lyapunov Functions (GLFs)
Let us consider (10) with an admissible $R$. Using results from [3, 4, 5], it can be shown that the following function: $V(x)=\max\mathcal{R}-\min\mathcal{R}$ [5], where =$\mathcal{R}=\{R_{1}(s,e),R_{2}(c_{1}),R_{3}(p,d),R_{4}(c_{2})\}$ is positive-definite with respect to the set of steady states and non-increasing for any choice of $R$. In addition, note that it can be written as $V={\widetilde{V}}\circ R$, where ${\widetilde{V}}$ is given as:
@@
Since $V$ is not assumed to be continuously differentiable, the gradient above is defined in terms of the upper Dini’s derivative as $\frac{\partial{{\widetilde{V}}}}{\partial r}(r)=[D_{e_{1}}^{+}{\widetilde{V}}(r),..,D_{e_{\nu}}^{+}{\widetilde{V}}(r)]$, where $\{e_{j}\}_{j=1}^{\nu}$ are the unit vectors and $D_{e_{j}}^{+}{\widetilde{V}}(r):=\limsup_{h\to 0^{+}}\tfrac{1}{h}({\widetilde{V}}(r+he_{j})-{\widetilde{V}}(r)).$
@@
(12) $\frac{\partial{\widetilde{V}}}{\partial r}K\Gamma r\;=\;\frac{\partial{\widetilde{V}}}{\partial r}\sum_{\ell=1}^{s}\bar{\rho}_{\ell}E_{i_{\ell}j_{\ell}}\Gamma r\;=\;\sum_{\ell=1}^{s}\bar{\rho}_{\ell}\overbrace{(e_{j_{\ell}}\gamma_{i_{\ell}}^{T})}^{Q_{\ell}}r\;=\;\sum_{\ell=1}^{s}\bar{\rho}_{\ell}Q_{\ell}r\leq 0\,,$
where $e_{j}$ is a member of the standard basis of $\mathbb{R}^{\nu}$, and $\gamma_{i}$ is the $i$th row of $\Gamma$. Since the coefficients $\bar{\rho}_{1},..,\bar{\rho}_{s}$ are arbitrary nonnegative numbers, this motivates the following definition:
@@
Let ${\widetilde{V}}:\mathbb{R}^{\nu}\to\mathbb{R}_{\geq 0}$ be a locally-Lipschitz function, and let $Q_{1},..,Q_{s}\in\mathbb{R}^{\nu\times\nu}$. Then, we say that ${\widetilde{V}}$ is a common Lyapunov function for the set of linear systems
$\dot{r}_{1}=Q_{1}r,\;\ldots\;,\dot{r}_{s}=Q_{s}r$
@@
A function ${\widetilde{V}}$ is an GLF for a given network iff ${\widetilde{V}}$ is a common Lyapunov function for the set of linear systems $\dot{r}=Q_{1}r,...,\dot{r}=Q_{s}r$.
@@
It can be verified that the function ${\widetilde{V}}$ defined in (11) is indeed a common Lyapunov function for the set of rank-one linear systems $\dot{r}=Q_{\ell}r,\ell=1,..,6$.

### v4

@@ 1 Introduction
Figure 1: The same BIN formalism $S+K\leftrightharpoons C\to P+K$ can represent processes as varied as: (a) an interaction between a catalyst (in the figure, an enzyme kinase) and its substrate, in which the two species can temporarily interact in a reversible fashion, sometimes resulting in the catalyst being released and the substrate being modified (“product”); (b) an interaction between an infective individual and a susceptible individual, in which the two individuals meet and either go their way without any new infection, or result in the susceptible individual becoming exposed to the disease; or (c) a predator/prey interaction in which the prey might not be harmed or the prey may end up injured or dead.
@@
We introduce the notion of structural contraction, and show how this property can be deduced from the conditions for stability introduced in [4, 5] and implemented in the computational package LEARN. In particular, we show that the existence of the rate-dependent Lyapunov functions proposed in [4, 5] for a given BIN implies structural nonexpansivity with respect to weighted $\ell_{\infty}$ norms, and strict contractivity on positive compact sets under additional but easily verifiable conditions. Furthermore, we show that such networks obey the notion of “contraction after a small overshoot and short transient” that was introduced in [63], and this fact allows us to show entrainment to periodic inputs. Therefore, the powerful tools of contraction theory can be applied to a rich class of BINs which can be characterized computationally via the MATLAB package LEARN (github.com/malirdwi/LEARN) or graphically [1].
The paper is organized as follows. We first review the main results. Section 2 discusses preliminaries including extensions of previous results that are relevant to our current analysis. Section 3 shows that nonexpansivity follows from the existence of an appropriate rate-dependent Lyapunov function. Section 4 establishes contraction over positive compact sets. Section 5 shows contraction after a small overshoot and short transient and entrainment to periodic inputs. Section 6 discusses biochemical examples in detail, and conclusions are discussed in Section 7.
@@ 1.1 Overview of Results
In this subsection, we offer an informal guide that summarizes the main results of the paper. They apply to BINs whose dynamics can be described by $\dot{x}=\Gamma R(x)$ where $\Gamma$ is the stoichiometry matrix and $R$ is the reaction rate. The only assumption needed is that the kinetics are monotone (refer to §2 for the exact assumptions).
@@
The results in [4, 5] provided functions that can be defined using the stoichiometry of the network without knowledge of the kinetics. We first define a rate-dependent Graphical Lyapunov Function (GLF) ${\widetilde{V}}(r)$, where $r$ has the dimension equal to the number of reactions, see Definition 2.1. It is characterized by the existence of a common Lyapunov function for a set of rank-one linear systems, see Theorem 2.4. This means that $V(x)={\widetilde{V}}(R(x))$ is a Lyapunov function which is always non-increasing along the dynamics, see Proposition 2.5. In addition, if $\widetilde{V}$} is invariant with respect to translation by $\Gamma$, then its existence is equivalent to the existence of a dual function $^{\widehat{V}}$ satisfying ${\widehat{V}}(\Gamma r)={\widetilde{V}}(r)$, see Theorem 2.12. Then, ${\widehat{V}}(x-{\bar{x}})$ is a Lyapunov function in the stoichiometric class compatible with a given steady state $\bar{x}$}, see Corollary 2.13.
@@
Our main result states that if an $\ell_{\infty}$ GLF ${\widetilde{V}}(r)=\|B\Gamma r\|_{\infty}$ exists for a network, then the norm of the difference between two trajectories that lie in the same stoichiometry class is non-increasing along time, see Theorem 3.4, thus establishing that the network is structurally nonexpansive with respect to the norm $\|.\|_{B}:z\mapsto\|Bz\|_{\infty}$. Furthermore, we provide an upper bound on the logarithmic norm of the Jacobian in terms of the partial derivatives of $R$, see Lemma 3.9.
@@
Given the satisfaction of these two verifiable conditions, we show that a network is structurally strictly contractive over positive compact sets with respect to the norm $\|.\|_{PB}:z\mapsto\|PBz\|_{\infty}$, where $P$ is an appropriately-defined diagonal matrix that might depend on the choice of the compact set, see Theorem 4.10.
@@
We then turn to the study of BINs where one or multiple parameters are periodic time-varying functions. We ask if all trajectories entrain to the same period. To that end, we use our result establishing contraction over positive compact sets to show when a given BIN is also contractive with short transients and small overshoots, see Theorem 5.5. Once this is known, the results from [63], allow us to conclude that trajectories of a periodic time-varying system that stay in the same stoichiometric class entrain to periodic inputs, see Theorem 5.7.
@@ 2.1 Notations
For any set $S$, $|S|$ denotes its cardinality. The set of $n$-tuples of positive numbers is denoted by $\mathbb{R}_{+}^{n}$, while the set of $n$-tuples of non-negative numbers is denoted by $\mathbb{R}_{\geq 0}^{n}$. The set of $n\times m$ matrices with real entries is denoted by $\mathbb{R}^{n\times m}$. For any subset $U\subset\mathbb{R}^{n\times m}$, its closure with respect to the standard Euclidean topology is denoted by $\bar{U}$. If $A\in\mathbb{R}^{n\times m}$ is a matrix, then ${\mathop{\rm Ker\,}}A$ is its kernel or null space, while ${\mathop{\rm Im\,}}A$ is its image space. For any function $V:\mathcal{X}\to\mathbb{R}$, ${\mathop{\rm Ker\,}}V:=\{x\in\mathcal{X}|V(x)=0\}$, and for a subset $\mathcal{Y}\subset\mathcal{X}$, ${\mathop{\rm Ker\,}}V|_{\mathcal{Y}}:=\{x\in\mathcal{Y}|V(x)=0\}$. For $x=[x_{1},\dots,x_{n}]^{T}\in\mathbb{R}^{n}$, $\|x\|_{\infty}:=\max_{i}|x_{i}|$. The notation $x\gg 0$ means that $x\in\mathbb{R}_{+}^{n}$, while $x\geq 0$ means that $x\in\mathbb{R}_{\geq 0}^{n}$. For a square matrix $A\in\mathbb{R}^{n\times n}$, $\sigma_{i}(A):=a_{ii}+\sum_{j\neq i}|a_{ij}|$, $i=1,\dots,n$, and $\mu_{\infty}(A):=\max_{i}\sigma_{i}(A)$. For a subspace $W\subset\mathbb{R}^{n}$, the quotient space $\mathbb{R}^{n}/W$ is the set of all sets of the form $\{x+y|y\in W\}$ where $x\in\mathbb{R}^{n}$.
@@ 2.2 Biological Interaction Networks
A Biological Interaction Network (BIN) $\mathscr{N}$ is specified as a pair $(\mathscr{S},\mathscr{R})$ where $\mathscr{S}:=\{X_{1},\dots,X_{n}\}$ is the set of symbols called species, and $\mathscr{R}:=\{{\rm\bf R}_{1},\dots,{\rm\bf R}_{\nu}\}$ is a set of symbols called reactions. Each reaction is associated to a formal rule of the following form:
${\rm\bf R}_{j}:~~\sum_{i=1}^{n}{\alpha_{ij}}X_{i}\longrightarrow\sum_{i=1}^{n}\beta_{ij}X_{i},$
where the nonnegative real numbers $\alpha_{ij},\beta_{ij}\geq 0$ are called the stoichiometric coefficients of the respective reaction. The $X_{i}$’s that appear with nonzero coefficients in the left-hand side are called the reactants of the reaction $j$, and the $X_{i}$’s that appear with nonzero coefficients in the right-hand side are called the products of the reaction $j$. In many applications the stoichiometric coefficients are integers, and one thinks of the $j$th reaction as saying that $\alpha_{{1j}}$ copies of $X_{1}$, …, $\alpha_{{nj}}$ copies of $X_{n}$ combine to produce $\beta_{{1j}}$ copies of $X_{1}$, …, $\beta_{{nj}}$ copies of $X_{n}$. The net “loss” (if negative) or “gain” (if positive) of the $i$th species in the $j$th reaction is $\gamma_{ij}=\beta_{ij}-\alpha_{ij}$. These numbers can be collected into a matrix $\Gamma\in\mathbb{R}^{n\times\nu}$ defined entry-wise as $\Gamma]_{ij}=\gamma_{ij}$. The matrix $\Gamma$ is customarily called the stoichiometry matrix.
@@
In order to quantify elements of the network, the species $X_{1},\dots,X_{n}$ are assigned non-negative numbers known as the concentrations (representing the density of elements in a certain region of space) $x_{1},\dots,x_{n}$, while reactions ${\rm\bf R}_{1},\dots,{\rm\bf R}_{\nu}$ are assigned rates $R_{j}:\mathbb{R}_{\geq 0}^{n}\to\mathbb{R}_{\geq 0},j=1,\dots,\nu$. These rates describe “how fast” each reaction can occur, depending on the concentration of its reactants. The most common form of reaction rates is known as mass-action kinetics, which are those of the form $R_{j}(x)={k_{j}}\prod_{i=1,\alpha_{ij}>0}^{n}{x_{i}}^{\alpha_{ij}}$, where $k_{j}>0$ is known as the kinetic constant. However, here we do not need to restrict to a specific form of kinetics. Instead we only assume that each such $R_{j}$ is monotone with respect to its reactants. More precisely, a reaction rate $R$ is admissible if it satisfies the following general properties:
@@
    each reaction varies smoothly with respects to its reactants, i.e. $R(x)$ is $\mathscr{C}^{1}$;
@@
    if a reactant increases, then the reaction rate increases, i.e. ${\partial R_{j}}/{\partial x_{i}}(x)\geq 0$ if $\alpha_{ij}>0$ and ${\partial R_{j}}/{\partial x_{i}}(x)\equiv 0$ if $\alpha_{ij}=0$. Furthermore, the aforementioned inequality is strict whenever the reactants are strictly positive.
@@
Another equivalent formulation is obtained by decomposing $R$ into a conic sum of indicator matrices. To that end, let $E_{ji}\in\{0,1\}^{\nu\times n}$ be an indicator matrix, i.e., $E_{ji}$ is zero everywhere except for the $(j,i)$th entry where it is equal to 1. In addition, let the elements of $\mathcal{P}_{\mathscr{N}}$ be indexed as $(i_{1},j_{1}),\dots,(i_{s},j_{s})$. Therefore, we can write
(2) $\mathcal{K}_{\mathscr{N}}=\{K\in\mathbb{R}^{\nu\times n}|K=\sum_{\ell=1}^{s}\bar{\rho}_{\ell}E_{j_{\ell}i_{\ell}},~\mbox{for some}~\bar{\rho}_{1},\dots,\bar{\rho}_{s}>0\}.$
Hence, the admissibility of $R$ implies the existence of continuous functions $\rho_{1}(x),\dots,\rho_{s}(x)\geq 0$ such that
@@
The time evolution of the concentration vector $x=[x_{1},\dots,x_{n}]^{T}$ as a function of time is by definition described by the following vector Ordinary Differential Equation (ODE):
@@
A non-zero vector $w\in\mathbb{R}_{\geq 0}^{n}/\{0\}$ is called a conservation law if $w^{T}\Gamma=0$. If there is a $w$ that is strictly positive then the corresponding network is called conservative. Note that this implies that $w^{T}\dot{x}(t)=0$ for all $t$, so $w^{T}x$ is a conserved quantity that stays invariant, i.e., $w^{T}x(t)\equiv w^{T}x(0)$. If such a $w$ exists, all trajectories will be confined to translates of the proper subspace ${\mathop{\rm Im\,}}\Gamma$ as can be seen by integrating (4) to get $x(t)=x_{0}+\Gamma\int_{0}^{t}R(x(\tau))\,d\tau$. Therefore, $x(t)\in\mathscr{C}_{x_{0}}$ for all $t\geq 0$ where
@@
The forward-invariant subset $\mathscr{C}_{x_{0}}$ is a closed and convex set known as the stoichiometric class corresponding to $x_{0}$. A stoichiometric class is proper if it intersects the open positive orthant.
@@
Conversely, suppose that $x_{0}$ is arbitrary and that $\xi$ is any solution of Equation (7), so that $\xi(t)=\int_{0}^{t}R(x_{0}+\Gamma\xi(\tau))\,d\tau$. Then we may define
@@
$\dot{x}(t)\;=\;\Gamma\dot{\xi}(t)\;=\;\Gamma R(x_{0}+\Gamma\xi({t}))\;=\;\Gamma R(x({t})).$
@@ 2.3 Graphical Lyapunov Functions (GLFs)
Let us consider (10) with an admissible $R$. Using results from [3, 4, 5], it can be shown that the following function: $V(x)=\max\mathcal{R}-\min\mathcal{R}$ [5], where $\mathcal{R}=\{R_{1}(s,e),R_{2}(c_{1}),R_{3}(p,d),R_{4}(c_{2})\}$ is positive-definite with respect to the set of steady states and non-increasing for any choice of $R$. In addition, note that it can be written as $V={\widetilde{V}}\circ R$, where $\widetilde{V}$} is given as:
@@
Since $V$ is not assumed to be continuously differentiable, the gradient above is defined in terms of the upper Dini’s derivative as $\frac{\partial{{\widetilde{V}}}}{\partial r}(r)=[D_{e_{1}}^{+}{\widetilde{V}}(r),\dots,D_{e_{\nu}}^{+}{\widetilde{V}}(r)]$, where $\{e_{j}\}_{j=1}^{\nu}$ are the unit vectors and $D_{e_{j}}^{+}{\widetilde{V}}(r):=\limsup_{h\to 0^{+}}\tfrac{1}{h}({\widetilde{V}}(r+he_{j})-{\widetilde{V}}(r)).$
@@
(12) $\frac{\partial{\widetilde{V}}}{\partial r}K\Gamma r\;=\;\frac{\partial{\widetilde{V}}}{\partial r}\sum_{\ell=1}^{s}\bar{\rho}_{\ell}E_{j_{\ell}i_{\ell}}\Gamma r\;=\;\sum_{\ell=1}^{s}\bar{\rho}_{\ell}\overbrace{(e_{j_{\ell}}\gamma_{i_{\ell}}^{T})}^{Q_{\ell}}r\;=\;\sum_{\ell=1}^{s}\bar{\rho}_{\ell}Q_{\ell}r\leq 0\,$,
where $e_{j}$ is a member of the standard basis of $\mathbb{R}^{\nu}$, and $\gamma_{i}$ is the $i$th row of $\Gamma$. Since the coefficients $\bar{\rho}_{1},\dots,\bar{\rho}_{s}$ are arbitrary nonnegative numbers, this motivates the following definition:
@@
Let ${\widetilde{V}}:\mathbb{R}^{\nu}\to\mathbb{R}_{\geq 0}$ be a locally-Lipschitz function, and let $Q_{1},\dots,Q_{s}\in\mathbb{R}^{\nu\times\nu}$. Then, we say that $\widetilde{V}$} is a common Lyapunov function for the set of linear systems
${\dot{r}}=Q_{1}r,{\ldots},\dot{r}=Q_{s}r$
@@
A function $\widetilde{V}$} is a GLF for a given network iff $\widetilde{V}$} is a common Lyapunov function for the set of linear systems $\dot{r}=Q_{1}r,\dots,\dot{r}=Q_{s}r$.
@@
It can be verified that the function $\widetilde{V}$} defined in (11) is indeed a common Lyapunov function for the set of rank-one linear systems $\dot{r}=Q_{\ell}r,\ell=1,\dots,6$.